What is the velocity of a 0.03 kg bullet that has a KE of 3750 J when fired from a rifle?And construct the free-body diagrams?


1
Expert's answer
2022-06-17T08:14:59-0400

The kinetic energy is given by

"E_k=\frac{mv^2}{2}"

The velocity of bullet

"v=\sqrt{2E_k/m}=\sqrt{2*3750/0.03}=500\:\rm m/s"

The free-body diagram is as follows